COMMON TOPICS OF INTEREST IN GAS AND CONDENSED PHASE 2D SPECTROSCOPIES: A BRIEF OVERVIEW

\underline{MARTIN T. ZANNI}, DEPARTMENT OF CHEMISTRY, UNIVERSITY OF WISCONSIN AT MADISON, MADISON, WI 53706.

This talk will give a brief overview of topics common to multidimensional spectroscopies.Exciting new advances have recently been taking place in the gas and condensed phase communities that are linked by common experimental methodologies and theoretical underpinnings. One example is double resonance spectroscopies that are now being performed in gases and liquids to expose conformationally sensitive couplings. Another common topic is how vibrational relaxation might be used to enhance structural information. The goal of this minisymposium is to help transfer ideas and foster collaborations between the two communities.